# What Do Privacy Advertisements Communicate to Consumers?

### v1

@@ Introduction
[Interviewer asks different variations of the question based on participants’ assigned treatment group and screening survey results. For Apple and Samsung groups we use the term “smartphone”; for WhatsApp we use “instant messaging”; fro Twitter we use “social media”; and for DuckDuckGo we use “browser and search engine”. For all smartphone groups we use the term “device”; for apps groups we use “service.”]
@@
Thank you, you can now stop screen sharing, Please return to Twitter and log out of the account.
@@
Q.11: On a scale of 1 - 5, with 1 being Poor and 5 being Excellent, how well do you think the (ad — game) communicated that message? {adjustwidth}0.5cm [Prompt] Why did you give it that rating?
[Q12A is for the ad campaign participants, 12B for the game participants.]
@@
Q.15: Relating back to the ad, What actions do you think the advertisement was suggesting you could take?
@@
Q.17: On a scale of 1 not at all motivated to 5, extremely motivated, how would you rank your motivation to (perform the steps — go to the game-level setting) shown in the (advertisement — game)?
Q.18: On a scale of 1 not at all confident to 5, fully confident, how would you rank your confidence in being able to (perform the steps — go to the game-level setting) shown in the (advertisement — game)?
@@
[Link phone to screen share into Zoom - Allow for 3 minute max for both activty (1) and (2) combined, do not reveal time to participants]
@@
[Watch for activity time, do not exceed 3 minutes max]
@@
Q.21: Did the (ad — game) motivate you to consider switching from your current (Apple & Samsung - smartphone — WhatsApp - instant messaging — Twitter - social media — DuckDuckGo - browser and search engine) to the (ad — game)’s (smartphone — instant messaging — social media — browser)? {adjustwidth}0.5cm [Prompt] Why?
@@
[Twitter skips Q23, interview completed.]
@@ A.3. Ads Screens
(c) A screen showing medical web page using DuckDuckGo showing trackers blocked.
@@
[…]
apple-01
(a) Data auction.
[…]
apple-02
(b) The protagonist using the “Ask App Not To Track” feature.
[…]
apple-03
(c) The protagonist using the “Protect Mail Activity” feature.
[…]
apple-04
(d) The auction participants disappearing into smoke after the protagonist used Apple’s privacy features.
Figure 10. Examples of Apple’s ad screens.
@@
[…]
samsung-01
(a) Samsung’s privacy dashboard showing which apps accessing which permission (e.g. camera, microphone, and location).
[…]
samsung-02
(b) Protagonist turned off precise location for the Weather app, portrayed by a yellow umbrella.
[…]
samsung-03
(c) Protagonist turned off precise location for the Weather app.
[…]
samsung-04
(d) Protagonist uses location access to control when to allow an app access their location.
Figure 12. Examples of Samsung’s ad screens.
[…]
whatsapp-01
(a) The post office.
[…]
whatsapp-02
(b) The post office officer uses pigeons as a message delivery method.
[…]
whatsapp-03
[…]
whatsapp-04
(d) The ad states that with WhatsApp, your messages will not be one of them (the encrypted messages).
Figure 13. Examples of WhatsApp’s ad screens.
@@ A.4. Participant Overview
Table 2. Participant Overview. In the “User Type” column, U represent users, and N represents non-users, W represents WhatsApp, and D represents DuckDuckGo. The “Tech” column indicates whether participants indicated in the screening survey that they had any experience learning or working in information technology or a related field.
@@ A.6. Codebook
Table 3. Codebook for Q.1 about participants’ preferred product. Note that the participant preferred products can be different from the product of the study’s group the participant was assigned to.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
engine-ddg (1) The DuckDuckGo search engine “Sometimes I’ll use DuckDuckGo or Bing” (WD2, DuckDuckGo)
engine-bing (1) The Bing search engine “Sometimes I’ll use DuckDuckGo or Bing” (WD2, DuckDuckGo)
@@
Table 4. Codebook for Q.1A about reasons for participants’ preferred product. Note that the participant preferred products can be different from the product of the study’s group the participant was assigned to.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 5. Codebook for Q.5 & Q.14 about participants’ impression about the advertising organization. The codes used for both pre-exposure (Q.5) or post-exposure (Q.14) to the campaign, except codes prefixed with “-plus” are applicable only for post-exposure.
@@
Table 6. Codebook for Q.5 about reasons for participants’ positive impression about the advertising organization pre-exposure to the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
reliable-product (1) The product is reliable “I know whenever SnapChat is down or something, they’ll always have people go straight to Twitter to talk about it” (T6)
@@
Table 7. Codebook for Q.5 about reasons for participants’ negative impression about the advertising organization pre-exposure to the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 8. Codebook for Q.5 about reasons for participants’ neutral impression about the advertising organization pre-exposure to the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 9. Codebook for Q.14 about change in participants’ impression about the advertising organization post-exposure to the campaign.
@@
Table 10. Codebook for Q.14 about reasons for change in participants’ impression about the advertising organization post-exposure to the campaign. Note that the number of occurrence of each code here represent the aggregate number where this code occurred in any combination of impressions change between pre- and post-exposure to the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 11. Codebook for Q.6 about participants’ prior-exposure to the campaign.
@@
Table 12. Codebook for Q.23 about the effect of re-exposure to the ad campaign on participants.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
more-questions (2) The participant has more questions of the ad “So does that mean the auctioneer is Apple. and by the consumer saying, Ask app not to track. Is that just us preventing Apple legally from being able to sell that data?” (A5)
@@
boring (1) The participant would feel boring if they keep watching it over and over “if I would want to watch this advertisement like a 100 times. I’m sure I’ll get sick of it.” (WD1, WhatsApp)
@@
Table 13. Codebook for Q.7, Q.8, & Q.9 about participants’ privacy awareness.
@@
Table 14. Codebook for Q.7, Q.8, & Q.9 about types of participants’ privacy awareness.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 15. Codebook for Q.7, Q.8, & Q.9 about types of privacy issues the participant is aware of.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 16. Codebook for Q.7, Q.8, & Q.9 about types of privacy features the participant is aware of.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 17. Codebook for Q.9 about participants’ perceived intent of the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 18. Codebook for Q.13 about participants’ prior awareness of the the advertised feature.
@@
Table 19. Codebook for Q.13B about how participants came across the advertised feature.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 20. Codebook for Q.10 about participants’ perception of the metaphors shown in the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 21. Codebook for Section 6 & 7 about the actionability task evaluation criteria.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 22. Codebook for Q.21 about participants’ (non-users of the product) motivation to switch to the advertised product post-exposure to the campaign.
@@
Table 23. Codebook for Q.21 about reasons for participants (non-users of the product) motivation to switch to the advertised product post-exposure to the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 24. Codebook for Q.21 about reasons for participants (non-users of the product) motivation to not-switch to the advertised product post-exposure to the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
same-company (1) The advertised product is owned by same company of the current used product (e.g. Meta) “they’re from the same company, and I really only use like Facebook messenger for just to contact friends with [too] personal stuff … and it’s owned by the same company” (WD2, WhatsApp)
@@
Table 25. Codebook for Q.21 about reasons for participants (non-users of the product) motivation to maybe-switch to the advertised product post-exposure to the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 26. Codebook for Q.22 about participants (users of the product) motivation to continue using the advertised product post-exposure to the campaign.
@@
Table 27. Codebook for Q.22 about reasons for participants (users of the product) motivation to continue using the advertised product post-exposure to the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 28. Codebook for Q.22 about reasons for participants (users of the product) to have no-impact on motivation to continue using the advertised product post-exposure to the campaign.
@@
Code (# occurrence) [non-exclusive] Code Definition Example
@@
Table 29. Codebook for Q.22 about reasons for participants (users of the product) to have neutral motivation to continue using the advertised product post-exposure.
@@
Code (# occurrence) [non-exclusive] Code Definition Example

### v2

@@ Introduction
[Interviewer asks different variations of the question based on participants’ assigned treatment group and screening survey results. For Apple and Samsung groups we use the term “smartphone;” for WhatsApp we use “instant messaging;” for Twitter we use “social media;” and for DuckDuckGo we use “browser and search engine.” For all smartphone groups we use the term “device;” for apps groups we use “service.”]
@@
[Time is up] Thank you, you can now stop screen sharing, Please return to Twitter and log out of the account.
@@
Q.11: On a scale of 1 - 5, with 1 being “Poor” and 5 being “Excellent,” how well do you think the (ad — game) communicated that message? {adjustwidth}0.5cm [Prompt] Why did you give it that rating?
[Q.12A is for the ad campaign participants, Q.12B for the game participants.]
@@
Relating back to the ad,
Q.15: What actions do you think the advertisement was suggesting you could take?
@@
Q.17: On a scale of 1 “Not at all motivated” to 5 “Extremely motivated,” how would you rank your motivation to (perform the steps — go to the game-level setting) shown in the (advertisement — game)?
Q.18: On a scale of 1 “Not at all confident” to 5 “Fully confident,” how would you rank your confidence in being able to (perform the steps — go to the game-level setting) shown in the (advertisement — game)?
@@
[Link phone to screen share into Zoom - 3 minutes max for both activity (1) and (2) combined, do not reveal time to participants]
@@
[Watch for activity time, 3 minutes max]
@@
Q.21: Did the (ad — game) motivate you to consider switching from your current (Apple & Samsung - smartphone — WhatsApp - instant messaging — Twitter - social media — DuckDuckGo - browser and search engine) to the (ad — game)’s (smartphone — instant messaging — social media — browser and search engine)? {adjustwidth}0.5cm [Prompt] Why?
@@
[Twitter skips Q.23, interview completed.]
@@ A.3. Ads Screens
[…]
apple-01
(a) Data auction.
[…]
apple-02
(b) The protagonist uses the “Ask App Not To Track” feature.
[…]
apple-03
(c) The protagonist uses the “Protect Mail Activity” feature.
[…]
apple-04
(d) The auction participants disappeared into smoke after the protagonist used Apple’s privacy features.
Figure 9. Examples of Apple’s ad screens.
[…]
samsung-01
(a) Samsung’s privacy dashboard showing which apps accessing which permission (e.g., camera, microphone, and location).
[…]
samsung-02
(b) The protagonist turned off the precise location for the Weather app, portrayed by a yellow umbrella.
[…]
samsung-03
(c) The protagonist turned off the precise location for the Weather app.
[…]
samsung-04
(d) The protagonist uses location access to control when to allow an app access their location.
Figure 10. Examples of Samsung’s ad screens.
[…]
whatsapp-01
(a) The post office.
[…]
whatsapp-02
(b) The post office clerk uses pigeons as a message delivery method.
[…]
whatsapp-03
[…]
whatsapp-04
(d) The ad states that with WhatsApp, your messages (the encrypted messages) will not be one of them.
Figure 11. Examples of WhatsApp’s ad screens.
@@
(c) A screen showing a medical web page using DuckDuckGo showing trackers blocked.
@@ A.4. Participant Overview
Table 2. Participant Overview. In the “User Type” column, U represents users, N represents non-users, W represents WhatsApp, and D represents DuckDuckGo. The “Tech” column indicates whether participants indicated in the screening survey that they had any experience learning or working in information technology or a related field.
@@ A.6. Codebook
Table 3. The codebook for Q.1 about participants’ preferred products. Note that a participant’s preferred product can be different from the product of the treatment group the participant was assigned to.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
engine-ddg (1) The DuckDuckGo search engine “Sometimes I’$l$ use DuckDuckGo or Bing” (WD2, DuckDuckGo)
engine-bing (1) The Bing search engine “Sometimes I’$l$ use DuckDuckGo or Bing” (WD2, DuckDuckGo)
@@
Table 4. The codebook for Q.1A about the reasons for participants’ preferred products. Note that a participant’s preferred product can be different from the product of the treatment group the participant was assigned to.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 5. The codebook for Q.5 & Q.14 about participants’ impressions about the advertising organization. The codes used for both pre-exposure (Q.5) and post-exposure (Q.14) to the campaign, except codes prefixed with “-plus” are applicable only for post-exposure.
@@
Table 6. The codebook for Q.5 about the reasons for participants’ positive impressions about the advertising organization pre-exposure to the campaign.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
reliable-product (1) The product is reliable “I know whenever SnapChat is down or something, they’$l$ always have people go straight to Twitter to talk about it” (T6)
@@
Table 7. The codebook for Q.5 about the reasons for participants’ negative impressions about the advertising organization pre-exposure to the campaign.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 8. The codebook for Q.5 about the reasons for participants’ neutral impressions about the advertising organization pre-exposure to the campaign.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 9. The codebook for Q.14 about the change in participants’ impressions about the advertising organization post-exposure to the campaign.
@@
Table 10. The codebook for Q.14 about the reasons for the change in participants’ impressions about the advertising organization post-exposure to the campaign. Note that the frequency of each code here represents the aggregate number where this code occurred in any combination of impressions change between pre- and post-exposure to the campaign.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 11. The codebook for Q.6 about participants’ prior-exposure to the campaign.
@@
Table 12. The codebook for Q.23 about the effect of re-exposure to the ad campaign on participants.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
more-questions (2) The participant has more questions of the ad “So does that mean the auctioneer is Apple? and by the consumer saying, Ask app not to track. Is that just us preventing Apple legally from being able to sell that data?” (A5)
@@
boring (1) The participant would feel boring if they keep watching it over and over “if I would want to watch this advertisement like a 100 times. I’m sure I’$l$ get sick of it.” (WD1, WhatsApp)
@@
Table 13. The codebook for Q.7, Q.8, & Q.9 about participants’ privacy awareness.
@@
Table 14. The codebook for Q.7, Q.8, & Q.9 about the types of participants’ privacy awareness.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 15. The codebook for Q.7, Q.8, & Q.9 about the types of privacy issues participants are aware of.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 16. The codebook for Q.7, Q.8, & Q.9 about the types of privacy features participants are aware of.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 17. The codebook for Q.9 about participants’ perceived intent of the campaign.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 18. The codebook for Q.13 about participants’ prior awareness of the the advertised features.
@@
Table 19. The codebook for Q.13B about how participants came across the advertised features.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 20. The codebook for Q.10 about participants’ perception of the metaphors shown in the campaigns.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 21. The codebook for Section 6 & 7 about the actionability task evaluation criteria.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 22. The codebook for Q.21 about participants’ (non-users of the product) motivation to switch to the advertised product post-exposure to the campaigns.
@@
Table 23. The codebook for Q.21 about reasons for participants’ (non-users of the product) motivations to switch to the advertised products post-exposure to the campaigns.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 24. The codebook for Q.21 about the reasons for participants (non-users of the product) motivations to not-switch to the advertised products post-exposure to the campaigns.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
same-company (1) The advertised product is owned by same company of the current used product (e.g., Meta) “they’re from the same company, and I really only use like Facebook messenger for just to contact friends with [too] personal stuff … and it’s owned by the same company” (WD2, WhatsApp)
@@
Table 25. The codebook for Q.21 about the reasons for participants’ (non-users of the product) motivations to maybe-switch to the advertised products post-exposure to the campaigns.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 26. The codebook for Q.22 about participants’ (users of the product) motivations to continue using the advertised products post-exposure to the campaigns.
@@
Table 27. The codebook for Q.22 about the reasons for participants’ (users of the product) motivations to continue using the advertised products post-exposure to the campaigns.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 28. The codebook for Q.22 about the reasons for participants’ (users of the product) no-impact on motivations to continue using the advertised products post-exposure to the campaigns.
@@
Code (# frequency) [non-exclusive] Code Definition Example
@@
Table 29. The codebook for Q.22 about the reasons for participants’ (users of the product) neutral motivations to continue using the advertised products post-exposure to the campaign.
@@
Code (# frequency) [non-exclusive] Code Definition Example